OpenStudy (anonymous):

Is this quadratic vertex form?? y+3=1(x+1)^2

OpenStudy (anonymous):

\[ (y+3)=4 \times \frac 1 4 \times (x+1)^2 \] The vertex is at (-1,3) and yes I think this is what they call vertex ...

OpenStudy (anonymous):

EDIT: (-1,-3)